Show that the uncertainty in the measurement of a physical quantity "f" is equal to the following expression:

"\u0394f=\u221a(\u27e8f^2 \u27e9-\u27e8f\u27e9^2 )"


1
Expert's answer
2021-07-20T15:18:40-0400

Assume

Displacement operator

We know that displacement operator

"∆f=\frac{\hbar}{mw}(n+\frac{1}{2})"

<f>=0

"<f^2>=\frac {\hbar}{mw}(n+\frac{1}{2})"

"∆f=\sqrt{<f^2>-<f>^2}"


Put value in eqution

"∆f=\sqrt{[\frac{\hbar}{mw}(n+\frac{1}{2})]^2-0}"

"∆f=\frac{\hbar}{mw}(n+\frac{1}{2})"

Left hand side = right hand side

Hence proved